OpenStudy (anonymous):

solve the system of equations 2x-5y=6 8x+3y=63

OpenStudy (anonymous):

x=333/46 and y=39/23

OpenStudy (anonymous):

7.2.? 1.6.?

OpenStudy (anonymous):

about yeah

OpenStudy (anonymous):

I don't think so it cant have a decimal

OpenStudy (anonymous):

use fractions then

OpenStudy (anonymous):

I don't know how to convert them into fractions.